Zeronknight (zeronknight):

Dawn completed 8 chores in 5 hours. What is her unit rate of chores per hour? * 1.6 to 1 8 to 1 0.625 to 1 5 to 1

Zeronknight (zeronknight):

@sammixboo

sammixboo (sammixboo):

What do you believe it is?

sammixboo (sammixboo):

Remember to set up a proportion \(\tt \dfrac{8}{5}~=~\dfrac{?}{1}\)

Zeronknight (zeronknight):

Ok, so 8 / 5 = ??

Zeronknight (zeronknight):

It's A? @sammixboo

sammixboo (sammixboo):

Yes :)